# The Economics of Information Pollution in the Age of AI: A General Equilibrium Approach to Welfare, Measurement, and Policy

### v1

@@ Abstract
The advent of Large Language Models (LLMs) represents a fundamental shock to the economics of information production. By asymmetrically collapsing the marginal cost of generating low-quality, synthetic content while leaving high-quality production costly, AI systematically incentivizes information pollution. This paper develops a general equilibrium framework to analyze this challenge. We model the strategic interactions among a monopolistic platform, profit-maximizing producers, and utility-maximizing consumers in a three-stage game. The core of our model is a production technology with differential elasticities of substitution ($\sigma_{L}>1>\sigma_{H}$), which formalizes the insight that AI is a substitute for labor in low-quality production but a complement in high-quality creation. We prove the existence of a unique ”Polluted Information Equilibrium” and demonstrate its inefficiency, which is driven by a threefold market failure: a production externality, a platform governance failure, and an information commons externality. Methodologically, we derive a theoretically-grounded Information Pollution Index (IPI) with endogenous welfare weights to measure ecosystem health. From a policy perspective, we show that a first-best outcome requires a portfolio of instruments targeting each failure. Finally, considering the challenges of deep uncertainty, we advocate for an adaptive governance framework where policy instruments are dynamically adjusted based on real-time IPI readings, offering a robust blueprint for regulating information markets in the age of AI.
Introduction
The rise of large language models (LLMs) marks a significant turning point in information economics, revolutionizing content production and profoundly impacting market efficiency, social welfare, and institutional stability. Unlike traditional information asymmetry, the challenge posed by LLMs lies in the near-zero marginal cost of producing low-quality content that appears plausible but lacks factual support or originality. This undermines traditional quality signals, as the production cost of high-quality information (which requires expertise and rigorous verification) remains high. This asymmetry shifts the relative production costs of high- and low-quality content in information markets, fostering a proliferation of low-quality content. Industry reports indicate that API costs for generative AI have fallen by over $90\%$ in two years, while the annual economic losses from disinformation exceed \$100\\$100 billion, highlighting the profound impact of this shift.
Given this, the paper addresses four central research questions. First, how does the asymmetric nature of AI technology reshape the equilibrium in the market for information? Second, what are the precise mechanisms of market failure in this new equilibrium, and how can their welfare consequences be characterized? Third, can we construct a theoretically-grounded, operational metric to measure the health of the information ecosystem in real time? Fourth, what is the optimal design of a multi-instrument policy framework that can correct these novel market failures in an environment of deep technological uncertainty?
To answer these questions, we develop a general-equilibrium model of the information market. Our central theoretical contribution lies in formalizing the asymmetric technological shock through production functions with differential elasticities of substitution, capturing that AI capital is a gross substitute for human labor in low-quality content production ($\sigma_{L}>1$) but a gross complement in high-quality creation ($\sigma_{H}<1$).
The core contribution of this study lies in theoretically demonstrating the uniqueness and inefficiency of the ”polluted information equilibrium” and explicitly identifying its triple market failure. Methodologically, we propose the ”Information Pollution Index” (IPI), a theoretically-grounded, multi-dimensional composite index derived directly from our model to quantify ecosystem health. Policy-wise, we derive an optimal multi-instrument framework, showing that achieving the first-best outcome requires a combination of tools like Pigouvian taxes, content provenance standards, and platform fiduciary obligations to simultaneously address these market failures. Recognizing implementation challenges under uncertainty, we further propose an adaptive governance framework that leverages the IPI as a real-time feedback signal.
The remainder of the paper is organized as follows. Section II lays out the theoretical framework. Section III formally defines and characterizes the Polluted Information Equilibrium, including its existence and inefficiency. Section IV develops the Information Pollution Index (IPI) as a theory-based metric. Section V solves for the optimal policy portfolio. Section VI discusses robustness and the challenges of deep uncertainty, motivating an adaptive governance approach. Section VII concludes with policy implications and directions for future research.
Literature Review
The economic theoretical basis and impact mechanism of information pollution
The digital age, characterized by unprecedented information flow and technological advancements, has given rise to a novel and complex market failure: information pollution. While its theoretical foundations can be traced back to Akerlof’s (Akerlof 1970) ”lemon market” model, its manifestation in the contemporary digital environment presents unique, multi-dimensional complexities. On one hand, false information spreads rapidly through the topological structure of social networks (Vosoughi, Roy, and Aral 2018), exhibiting more significant network externalities (Allcott and Gentzkow 2017) and being continuously reinforced by cognitive biases (such as the truth illusion effect revealed by Pennycook &\& Rand (Pennycook and Rand 2019)). On the other hand, information pollution causes multidimensional economic distortions, such as (Allcott et al. 2020), (Durante, Pinotti, and Tesei 2019) and (Bursztyn et al. 2020).
If we consider it from the perspective of pollution, environmental economics has two important tools for dealing with pollution: Pigouvian taxes and tradable permits. A Pigovian tax is an economic tool designed to correct market failures caused by negative externalities. In recent years, some scholars have begun to explore the application of Pigovian taxes to digital platform governance. Dranove et al. (Dranove et al. 2014) proposed the idea of levying an ”attention tax” on online advertising to alleviate the problem of information overload. Cabral et al. (Cabral et al. 2021) further developed this idea and proposed a differentiated tax mechanism for low-quality content. Zuboff (Zuboff 2019) proposed the concept of ”information extraction tax” in The Age of Surveillance Capitalism, its theoretical framework is also inspiring for information pollution governance. Tradable permits, first proposed by Dales (Dales 1968) and based on the property rights theory of Coase (Coase 1960), that imposes a tradable hard cap (cap) on the total amount of pollutants that can be emitted by a particular industry or economy. Montgomery (Montgomery 1972) demonstrated that, under certain conditions, tradable permits can achieve pollution control objectives with minimal costs. Although there are some obstacles to shifting this perspective to information pollution, some scholars have put forward some relevant proposals. Parker &\& Van Alstyne (Parker and Van Alstyne 2021) proposed the transaction of information quality points between platforms to solve the problem of information asymmetry in the platform market. In a report written for the European Commission, Crémer et al. (Crémer, de Montjoye, and Schweitzer 2019) proposed the concept of ”content quotas”, similar to a tradable license mechanism.
The impact of AI on the information market: governance challenges and multidisciplinary integration research
There are significant challenges in applying traditional economic governance tools to information pollution governance. In the information field, these approaches are more likely to raise questions about fairness, effectiveness, and potential conflicts with the right to freedom of expression, and may lead to censorship or discriminatory law enforcement. In practical terms, it is also difficult for a tradable license system to clearly define the ”upper limit” of information pollution. Without a strong definition, it will be very difficult to implement such a system fairly and effectively, which may distort competition and consolidate the market dominance of large platforms. Existing understandings also lack the need for long-term analysis. All of this highlights the need for comprehensive impact assessments and flexible regulatory frameworks. This paper contributes to and bridges three major strands of literature. First, we extend the classic literature on information economics. We recast the “market for lemons” ((Akerlof 1970)) for the digital age, and we explore how AI undermines traditional signaling mechanisms ((Spence 1973)). Second, we contribute to the literature on platform economics and industrial organization ((Rochet and Tirole 2004); (Hagiu and Wright 2015)). Our model is the first to formally integrate a sophisticated production-side analysis of AI’s impact into a platform intermediation framework. Third, we build upon the burgeoning literature on the economics of AI ((Agrawal, Gans, and Goldfarb 2018); (Acemoglu and Restrepo 2019)). Unlike most research on the labor market, our focus is on its impact on the quality of output in one of society’s most critical markets: the information market.
Theoretical Framework
This section develops the core theoretical framework of the paper. We first specify the economic environment, agents, and types of information goods. Second, we characterize the asymmetric production technology for content, which serves as the fundamental driver of pollution. Finally, we formalize the strategic interactions among the platform, producers, and consumers as a three-stage Stackelberg game, outlining its equilibrium characteristics and the sources of market failure.
The Economic Environment
Agents.
Our economic system consists of four types of rational agents, each with well-defined objectives and constraints:
@@ Introduction
  • Content Producers: A continuum of profit-maximizing producers indexed by $i\in[0,1]$. Each producer $i$ is endowed with a heterogeneous productivity parameter $A_{j,i}$ and faces factor prices $(r,w)$, where $r$ is the rental rate of AI capital and $w$ is the wage rate for high-skilled human labor.
  • Monopolistic Platform: Acting as the intermediary, the platform wields market power and serves a dual role. As a market organizer, it sets the content moderation intensity $m\in[0,1]$. As an attention allocator, it determines the exposure weights for different content types via an algorithmic amplification vector $\bm{\gamma}=(\gamma_{H},\gamma_{L})$.
@@
  • Consumers: A representative group of consumers with heterogeneous verification costs $k_{i}$, drawn from a distribution $F(k)$ with support $0,\bar{k}]$. Consumers observe a noisy signal $s\in\{H,L\}$ of content quality and decide whether to pay $k_{i}$ to perfectly ascertain it.
  • Social Planner: A theoretical benchmark character whose objective is to maximize social welfare. The planner allows us to define the Pareto efficient allocation and quantify the welfare loss from market failures.
Information Goods.
Two fundamentally different types of information goods exist in the market:
  • High-Quality Content ($Q_{H}$): Information products characterized by factual accuracy, cognitive value that improves consumer decision-making, and production complexity requiring expertise and rigorous verification.
  • Low-Quality Content ($Q_{L}$): Information products that may be misleading or false, generate negative externalities by harming consumer welfare and social trust, and are simple to produce via templated or automated means.
This classification is based not on subjective judgment but on the objective impact of content on consumer decisions and social welfare.
Production Technology: The Dynamics of Pollution
CES Production Function Specification.

### v2

# The Economics of Information Pollution in the Age of AI: General Equilibrium, Welfare, and Policy Design

@@ Abstract
The rapid diffusion of generative artificial intelligence (AI) is triggering structural change in information markets, with impacts far more complex than simple ’technological progress’. This paper is the first to model AI as an asymmetric technological shock: in low-quality content production, AI acts as a substitute for labor ($\sigma_{L}>1$), whereas in high-quality creation, it functions merely as a complement to human expertise ($\sigma_{H}<1$). This technological asymmetry fundamentally alters the cost structure of information production, systematically favoring ’lemons’ over ’peaches’. By constructing a three-stage general equilibrium model, we prove this shock leads to an inefficient ’Polluted Information Equilibrium’ sustained by a triad of interacting market failures: (1) a Production Externality, as low-quality producers do not internalize ecological harm; (2) a Platform Governance Failure, as engagement-based revenue models misalign algorithms with social welfare; and (3) a Trust Commons Externality, as verification, a public good, is systematically under-provided. To address this, we design an adaptive governance framework, first deriving a theoretically-grounded Information Pollution Index (IPI) for real-time ecosystem monitoring. Second, we demonstrate that restoring efficiency requires a multi-instrument policy portfolio: a Pigouvian tax to correct the production externality, content provenance standards to address under-verification, and fiduciary duties to constrain platform behavior. Finally, through Agent-Based Model (ABM) validation, we find this portfolio generates superadditive welfare gains—the joint intervention is significantly more effective than any single policy tool. The core insight is that the welfare consequences of the AI revolution depend not on the technology itself, but on how we design market rules; without proper governance, technological progress can paradoxically reduce social welfare, a phenomenon we term the ’Paradox of AI Progress’.
1 Introduction
In 2025, an in-depth investigative report, the product of a senior journalist’s three-week effort, garners 100,000 views on social media. On the same day, a sensational but unverified ”scoop” generated by an AI in three seconds receives 10 million shares. This stark contrast reveals the fundamental challenge facing the information economy in the age of generative AI: when the marginal cost of producing convincing but unverified content approaches zero, how can truth compete with noise?
This is not merely a technical question; it is a core economic one. The rise of large language models (LLMs) like GPT-4 and Claude represents a structural shift in the information production function. However, existing research has largely focused on AI’s impact on the labor market, overlooking a more fundamental problem: how does AI alter the equilibrium distribution of information quality?
Consider two distinct content production scenarios:
  • Scenario A (Investigative Report): A journalist uses AI tools to accelerate data analysis and assist in literature review, but the core work—field interviews, fact-checking, and ethical judgment—still requires human expertise. Here, AI is a complement, not a substitute.
  • Scenario B (Content Farm): An operator uses AI to batch-generate plausible health advice, investment ”insider” tips, or political ”revelations”. No professional knowledge is needed, only prompt engineering. Here, AI is a substitute, almost entirely replacing human labor.
This asymmetry in production technology—what we term ”technology-biased information pollution”—is reshaping the digital information ecosystem. By some estimates, 90% of online content may be AI-generated or AI-assisted by 2026. If the majority of this is unverified, low-quality content, we face not just ”information overload,” but systemic ”information pollution”—a new class of market failure.
The central thesis of this paper is that generative AI is not a neutral productivity tool, but rather a biased technological shock that systematically lowers the production cost of ”lemons” (low-quality content) relative to ”peaches” (high-quality content). We argue that this technological bias leads to market failure through three mutually reinforcing mechanisms: at the micro-level, individual producers rationally choose to produce low-quality content due to its low cost and rapid, algorithm-driven rewards; at the meso-level, platform algorithms designed to maximize engagement amplify content based on virality rather than veracity; and at the macro-level, a systemic erosion of trust leads to ”verification fatigue” among consumers, further degrading the market’s ability to discern quality.
This ”pollution spiral,” if left unchecked, risks a ”tragedy of the digital commons,” where everyone is a victim, yet no single actor has the incentive to unilaterally change behavior.
This paper’s contributions are threefold: a theoretical innovation that is the first to model AI’s market impact as an asymmetric Constant Elasticity of Substitution (CES) production function, revealing a paradox of progress; a measurement tool that constructs a multi-dimensional Information Pollution Index (IPI) to provide a quantitative basis for policy intervention; and a policy design that proves the necessity of a multi-instrument portfolio, validated through computational simulation.
The remainder of this paper is organized as follows. Section 2 reviews the relevant literature and situates our contribution. Section 3 builds the theoretical model and characterizes the equilibrium. Section 4 derives the IPI and designs the optimal policy portfolio. Section 5 provides validation via an Agent-Based Model (ABM). Section 6 concludes and discusses policy implications.
Figure 1: Conceptual Framework of AI-Driven Information Pollution. The model proceeds from the core asymmetric shock ($\sigma_{L}>1>\sigma_{H}$) through a three-stage Stackelberg game, which leads to a triad of interacting market failures: a Production Externality, a Platform Governance Failure, and an Information Commons Externality. These failures produce a self-reinforcing, Pareto-inefficient ”Polluted Information Equilibrium.” We show this equilibrium necessitates a multi-instrument policy portfolio (Pigouvian Tax, Provenance Mandates, Fiduciary Duties) to restore efficiency.
2 Literature Review
2.1 Economic Foundations and Traditional Governance of Information
The novel market failure of information pollution builds on Akerlof’s foundational “market for lemons” model Akerlof (1970), but manifests in far more complex, multi-dimensional ways in today’s digital environment. From the standpoint of economic theory, environmental economics provides two classic policy instruments for managing pollution: Pigouvian taxes and tradable permits.
Posner and Weyl Posner and Weyl (2018), in their book Radical Markets, explicitly advocated for taxing online advertising. Cabral et al. Cabral et al. (2021) further developed this idea and proposed a differentiation mechanism for low-quality content. In Zuboff’s Zuboff (2019) book The Age of Surveillance Capitalism, her profound critique of technology companies’ massive acquisition of user data to predict and change human behaviour provides an important theoretical basis and justification for regulating data exploitation and information pollution through economic means (such as taxation).
The tradable permit system is a mature governance tool in environmental economics. The core of this mechanism is setting a strict cap on total pollution emissions (Cap) and allocating emission rights as a market-tradable asset. Its theoretical foundation stems from Coase’s Coase (1960) property rights theory and was first systematically proposed by Dales Dales (1968). Subsequently, Montgomery’s Montgomery (1972) research theoretically demonstrated that, under specific conditions, this mechanism can achieve predetermined emission reduction targets with the lowest total social cost. Although there are some obstacles to shifting this perspective to information pollution, some scholars have put forward some relevant proposals.
In addition, some scholars have proposed market-based solutions from the perspective of platform governance. For example, the research of Parker and Van Alstyne Parker and Van Alstyne (2010) inspired using reputation systems or information quality points to incentivise platforms to solve information asymmetry. Additionally, Crémer et al. Crémer et al. (2019) proposed mandatory data sharing and interoperability – essentially “access quotas” for core resources, similar to tradable Licences – to break monopolies and promote competition.
2.2 AI, Algorithmic Amplification, and Ecosystem Degradation
Beyond economics, scholars in information science and communication studies have emphasised the ecological nature of digital knowledge systems. Floridi Floridi (2010) conceptualises information environments as “infospheres”, arguing that pollution arises when the semantic integrity of these ecosystems is undermined. Bawden and Robinson Bawden and Robinson (2009) describe the resulting cognitive and social fatigue as a form of “information environmental degradation”. These perspectives reinforce the economic analogy: information pollution, like environmental pollution, depletes a scarce common resource – trust.
Recent literature underscores that the information environment has entered a phase where quantity and veracity are inversely correlated. False news spreads faster and farther than verified information Vosoughi et al. (2018), amplified by algorithmic incentives that optimise for engagement rather than truth Allcott and Gentzkow (2017); Allcott et al. (2019). This phenomenon is continuously reinforced by cognitive biases (such as the lazy thinking effect revealed by Pennycook and Rand Pennycook and Rand (2019)) and causes multidimensional economic distortions, such as those highlighted by Allcott et al. Allcott et al. (2020), Durante et al. Durante et al. (2019), and Bursztyn et al. Bursztyn et al. (2020).
Algorithmic amplification further intensifies these dynamics. As Tufekci Tufekci (2015) observes, recommendation systems can generate systemic harms even without malicious intent. Empirical evidence confirms that algorithmic curation narrows informational diversity Bakshy et al. (2015). From a systems perspective, Helbing Helbing (2018) frames these feedback loops as emergent properties of self-organising social systems, suggesting that agent-based modelling can illuminate how micro-level incentives produce macro-level information distortions.
This paper contributes to and bridges three major strands of literature. First, it extends the classic literature on information economics in the context of digital content, exploring how AI undermines traditional signalling mechanisms Spence (1973). The erosion of credibility parallels earlier discussions on asymmetric information and adverse selection, but now operates within algorithmically mediated attention markets. Second, it advances research in platform economics and industrial organisation Hagiu and Wright (2015); Rochet and Tirole (2003). By integrating AI-driven production asymmetry into a two-sided platform framework, we provide a formal treatment of how content generation technologies interact with intermediation and market power. Third, it contributes to the emerging field of AI economics Acemoglu and Restrepo (2019); Agrawal et al. (2018), shifting the focus from labour displacement to the quality of informational output in one of society’s most critical domains – the information market.
3 Theoretical Framework
3.1 Conceptual Foundation: Information as a Common-Pool Resource
3.1.1 The Credible Information Commons
We conceptualize the ecosystem of credible information as a common-pool resource (CPR) in the Ostromian sense, rivalrous in quality yet only imperfectly excludable in access. Let the stock of high-quality information at time $t$, denoted by $S_{t}$, represent the attention-weighted reservoir of credible content circulating across digital platforms. New content creation replenishes this reservoir through a regenerative flow $R_{t}$, while low-quality or misleading information generated by actors seeking visibility, clicks, or advertising rents degrades it at a pollution rate $P_{t}$. The dynamic evolution of the resource follows
$\dot{S}_{t}=R_{t}-P_{t}(S_{t},\,a_{t},\,\theta_{t}),$ (1)
where $a_{t}$ captures the aggregate extraction intensity of attention by content producers, and $\theta_{t}$ summarizes governance and verification institutions that regulate entry, provenance, and moderation. The credible-information commons is renewable but exhaustible: its replenishment depends on costly verification and sustained trust, while degradation arises from unpriced externalities of generative AI and algorithmic amplification.
Extraction in this domain refers to the appropriation of limited audience attention $A_{t}$ for private content visibility. A simple representation of allocation shares is $s_{i}=\frac{q_{i}A_{t}}{\sum_{j}q_{j}A_{t}}$. Pollution corresponds to the dissemination of low-credibility or synthetic material that reduces the expected information accuracy $\mathbb{E}[h_{i}]$ and undermines collective trust $T_{t}=f(S_{t})$. The marginal social cost of pollution, $\partial C^{soc}/\partial P_{t}$, is convex in $P_{t}$ because reputational erosion and verification fatigue accelerate once misinformation exceeds cognitive processing capacity.
To move from the tragedy to sustainable governance, Ostrom’s design principles can be adapted to digital information. Boundaries become traceable provenance and authentication of content and accounts, which can be parameterized by the verifiability intensity $m$. Congruence between rules and local conditions corresponds to aligning platform recommendation objectives $(\gamma_{H},\gamma_{L})$ with societal welfare by rewarding credible engagement over sheer volume. Collective choice arrangements imply participatory moderation and community fact-checking that internalize positive monitoring externalities. Monitoring and graduated sanctions suggest layered detection and proportionate penalties for repeated offenders, representable as increasing marginal cost of pollution $c_{P}(P_{t})$. Conflict-resolution mechanisms provide low-cost correction channels to restore truthful information without discouraging legitimate debate. Finally, nested governance combines platform-level rules with national and international regulatory oversight, which we capture by multi-layer institutional variables $\theta_{t}=(\theta^{plat},\theta^{nat},\theta^{intl})$.
3.1.2 The Tragedy of AI-Driven Pollution
Each producer chooses output $q_{i}$ and quality $h_{i}$ to maximize private returns,
$\pi_{i}=p(h_{i},q_{i})-c(h_{i},q_{i}),$ (2)
where $p(\cdot)$ is the monetized visibility price determined by platform algorithms. Because platforms typically reward engagement rather than veracity, producers internalize revenue from exposure but not the social cost of degraded credibility. In aggregate, rational behavior therefore implies $\partial P_{t}/\partial a_{t}>0$, generating a tragedy of the commons: collectively excessive attention extraction and quality dilution, even when each agent behaves optimally given existing rules. A social planner would internalize the marginal external damage $MEC(S_{t})$ imposed on the shared trust stock, but decentralized markets lack such a mechanism in the absence of governance or norms.
Linking these mechanisms, AI-enabled content generation and algorithmic amplification increase extraction $a_{t}$ and raise $P_{t}$ unless institutional strength $\theta_{t}$ and verifiability $m$ are sufficiently high. As $P_{t}$ grows, the trust stock $T_{t}=f(S_{t})$ depreciates faster, verification becomes costlier due to fatigue, and marginal damages become convex. These forces reinforce a high-pollution equilibrium.
3.1.3 Preview of the Three-Stage Game
The CPR perspective nests within the general-equilibrium model in Section LABEL:sec:model. Micro-level over-extraction of attention maps to the production externality; platform bias in $(\gamma_{H},\gamma_{L})$ represents governance failure; and the depreciation of trust, $\dot{T}_{t}<0$, corresponds to a trust commons externality.
The strategic environment unfolds in three stages. In Stage I (platform), recommendation parameters $(\gamma_{H},\gamma_{L})$ and verifiability intensity $m$ are chosen subject to retention and profitability objectives, under institutional constraints $\theta_{t}$. In Stage II (producers), agents choose $(q_{i},h_{i})$ given platform rules, internalizing exposure revenue but not social damage, thereby determining $a_{t}$ and $P_{t}$. In Stage III (consumers and trust), aggregate attention allocation $A_{t}$ and verification behavior update stock variables $(S_{t},T_{t})$ via $\dot{S}_{t}=R_{t}-P_{t}(\cdot)$ and a depreciation condition for trust when pollution dominates, feeding back into platform and producer incentives.
The institutional principles above act as constraints that can shift the equilibrium from a high-pollution Nash outcome toward a cooperative optimum. Subsequent empirical sections operationalize these dynamics through cross-country proxies of media governance and social-media usage, while simulation experiments evaluate policy instruments using the Information Pollution Index (IPI) as a quantitative measure of commons health.
3.2 Model Primitives and Environment
3.2.1 Agents and Objectives
@@
The economy features four classes of rational agents. First, there is a continuum of content producers indexed by $i\in[0,1]$. Producer $i$ has heterogeneous productivity $A_{j,i}$ across content types $j\in\{H,L\}$ and faces factor prices $(r,w)$, where $r$ is the rental cost of AI capital and $w$ is the wage for high-skill human labor. Given platform rules, each producer chooses output and quality to maximize profits.
Second, a monopolistic platform intermediates the market and holds market power. As a market organizer, it selects the intensity of moderation and provenance $m\in[0,1]$. As an attention allocator, it commits to an algorithmic amplification vector $\boldsymbol{\gamma}=(\gamma_{H},\gamma_{L})$ that maps content types into exposure weights. The platform’s objective is specified in the three-stage game below and trades off profit, retention, and compliance.
@@
Third, consumers have heterogeneous verification costs $k_{i}\sim F(k)$ with support $0,\bar{k}]$. Upon receiving a noisy signal of content quality, consumers decide whether to pay $k_{i}$ to perfectly verify quality, and then allocate attention and demand accordingly.
Fourth, a social planner provides a normative benchmark. The planner maximizes social welfare by internalizing the externalities that low-quality content imposes on the stock of trust and informational accuracy. This allows us to define the Pareto-efficient allocation and quantify welfare losses from market failure.
3.2.2 Information Goods: Quality Differentiation
There are two information goods reflecting different production difficulties, externalities, and social welfare impacts. High-quality content $Q_{H}$ is accurate and cognitively valuable, improving individual decision-making and typically requiring expertise and rigorous verification. Low-quality content $Q_{L}$ can be misleading or false, generates negative externalities by eroding consumer welfare and social trust, and is scalable via templated or automated generation. The classification is based on objective effects on decisions and welfare, not subjective judgments. Under algorithmic amplification and attention competition, AI and automation disproportionately reduce the marginal cost and raise the scalability of $Q_{L}$, creating asymmetric technology that underlies pollution pressure.
3.2.3 Timeline and Information Structure
The timing follows a three-stage Stackelberg game. In Stage 1 (platform choice), given institutional constraints and market conditions, the platform commits to moderation and provenance intensity $m$ and to an amplification vector $\boldsymbol{\gamma}=(\gamma_{H},\gamma_{L})$. These choices determine expected exposure and monetization for content types and shape downstream best responses.
In Stage 2 (producer choice), producers observe $(m,\boldsymbol{\gamma})$ and factor prices $(r,w)$, and choose output and quality $(q_{i},h_{i})$ based on their productivity vector $A_{j,i}$. Expected unit revenue is given by an exposure-price function $p(h_{i},q_{i};\boldsymbol{\gamma},m)$, while costs follow $c(h_{i},q_{i};r,w)$. Because platform rewards focus on engagement rather than veracity, individual producers do not internalize the social costs that low-quality content imposes on public trust and verification fatigue.
In Stage 3 (consumption and trust update), consumers observe a noisy signal $s\in\{H,L\}$ and form priors given the platform environment, then decide whether to pay $k_{i}$ to fully verify. The resulting attention allocation and demand determine exposure and effective consumption. At the end of the period, the stock of credible information and trust updates via a regeneration–pollution process, $S_{t+1}=S_{t}+R_{t}-P_{t}$, where $R_{t}$ and $P_{t}$ are jointly determined by the production and propagation of $Q_{H}$ and $Q_{L}$ and by the platform’s $(m,\boldsymbol{\gamma})$. Trust depreciation induced by pollution feeds back into next period’s verification incentives and platform trade-offs, creating dynamic interactions.
Regarding information, the platform’s $(m,\boldsymbol{\gamma})$ is publicly observed at Stage 1. In Stage 2, producers know platform rules, factor prices, and their own productivity, and they form rational expectations over consumer verification. In Stage 3, consumers observe the platform-provided signal and historical environment, including partially observable proxies for governance and media conditions, and then make optimal verification and consumption decisions. The equilibrium concept is subgame-perfect equilibrium: given platform commitments, producer and consumer best responses are mutually consistent, and the platform chooses $(m,\boldsymbol{\gamma})$ anticipating downstream reactions. This timeline and information structure underpin the subsequent equilibrium characterization, comparative statics, and welfare analysis.
Figure 2: Production Cost Asymmetry under AI Technological Progress (Simulation of Proposition 3.2). This figure illustrates the model’s core mechanism. Panel A: As AI capital cost ($r$) decreases (AI progresses), the unit cost of low-quality content ($c_{L}$, substitute, $\sigma_{L}>1$) falls disproportionately faster than that of high-quality content ($c_{H}$, complement, $\sigma_{H}<1$). Panel B: This is because the cost elasticity with respect to $r$ is higher for low-quality content ($|\partial\log c_{L}/\partial\log r|>|\partial\log c_{H}/\partial\log r|$). Panel C: The result is a widening absolute and relative cost gap, creating a systematic and growing economic advantage for ”lemons” as AI technology improves.
3.3 Production Technology: The Dynamics of Pollution
3.3.1 CES Production Function Specification.